# CoSpaDi: Compressing LLMS via Calibration-Guided Sparse Dictionary Learning

### v2

@@ A.1 Derivation of the optimal pair of basis and coefficient matrices
We are seeking the optimal pair $\left(\mathbf{B}^{\star},\mathbf{C}^{\star}\right)$ that minimizes the constrained problem in Eq. (3). First, we rewrite the objective $\mathcal{J}$ in its equivalent form:
@@ A.5 Low-rank and CoSpaDi Inference Complexity
With proposed truncation of 2-bits in mantissa we can omit term for storing indices of nonzero elements in $\mathbf{S}$ resulting in corrected compression ratio: ^γ^SD = 1 - d1k + s d2d1d2 The rank for the low-rank decomposition could be estimated from the compression ratio with the following equation: r=(1 - γLR) d1d2d1+ d2 For sparse dictionary based method we defined $\rho=k/s$ and, thus we can estimate both $k$ and $s$ in the following way: k=(1-^γSD) d1d2d1+d2ρ,  s=kρ. Under the same compression ratio for both SVD and CoSpaDi we obtain exactly the same complexity: O_LR=N r (d_1 + d_2) = N (1 - γ^LR) d_1 d_2 O_SD=N d_1 k + N s d_2 = N(d_1 k + d_2 kρ) = N k (d_1 + d2ρ) = N (1-^γSD) d1d2d1+d2ρ(d_1 + d2ρ)=N(1-^γ^SD) d_1d_2
@@
While, theoretical complexity is the same, in practice inference time depends on the sparsity structure ($K_{\text{active}}$), indexing overhead, and kernel efficiency, as can be observed from Figs. 4, 5 and   6.
@@
A.6 Comparison with Other Training-Free Methods
We also evaluate the performance of the proposed CoSpaDi relative to other training free methods, particularly structural pruning ones and ASVD Yuan et al. (2023) which is another data-aware SVD-based method. The results are provided in Table A.6.
Across LLaMA-3 8B, our method consistently outperforms structural pruning baselines at moderate budgets. At $\text{CR}\!\approx\!0.2$, CoSpaDi attains the best average accuracy (0.618 vs. 0.580 for ReplaceMe and 0.551 for LLM-Pruner) while also delivering the lowest perplexities, indicating balanced generative and discriminative quality. At $\text{CR}\!\approx\!0.3$, CoSpaDi widens the margin in average accuracy (0.537 vs. 0.469/0.405) and avoids the severe perplexity blow-ups observed for pruning methods. Even under aggressive compression ($\text{CR}\!\approx\!0.4$–0.5), CoSpaDi remains competitive or superior on most tasks and preserves far more stable perplexity profiles, whereas ReplaceMe and LLM-Pruner exhibit task collapse (near-zero LAMBADA and extreme PPL spikes). Overall, the results support dictionary–sparsity as a more robust data-free alternative to structural pruning at matched or tighter budgets, especially when both accuracy and perplexity must be maintained.
Table 5: Comparison of the proposed CoSpaDi method with other training-free methods including ASVD Yuan et al. (2023) and state-of-the-art structured pruning methods ReplaceMe Shopkhoev et al. (2025) and LLM-Pruner Ma et al. (2023) on Llama3 8B under different compression ratios. We report accuracy on different benchmarks as well as its average and perplexity. Best results are highlighted with bold
[…]
ASVD 0.2 0.5577 0.2666 0.0070 0.3237 0.2150 0.3050 0.2182 0.2347 0.2660 9.38E+04 9.89E+05
ReplaceMe 0.22 0.7307 0.6572 0.4211 0.6587 0.4369 0.8640 0.3541 0.5167 0.5799 3.37E+01 2.01E+01
LLM-Pruner 0.7546 0.6746 0.5096 0.6208 0.3660 0.8780 0.3512 0.2495 0.5505 1.60E+01 1.05E+01
CoSpaDi 0.2 0.7519 0.6649 0.7380 0.6654 0.4155 0.8950 0.3818 0.4282 0.6176 1.97E+01 4.27E+00
ASVD 0.3 0.5359 0.2595 0.0000 0.2694 0.2142 0.2190 0.2115 0.2349 0.2431 3.43E+05 1.35E+07
ReplaceMe 0.31 0.6659 0.5382 0.2401 0.5067 0.3788 0.7730 0.3397 0.3058 0.4685 6.67E+01 1.33E+02
LLM-Pruner 0.6725 0.4509 0.2090 0.4541 0.2875 0.6340 0.3014 0.2292 0.4048 3.79E+01 2.21E+02
CoSpaDi 0.3 0.7051 0.5620 0.6128 0.5421 0.3353 0.8570 0.3617 0.3224 0.5373 4.46E+01 9.18E+00
ASVD 0.4 0.5365 0.2575 0.0000 0.2504 0.2287 0.2080 0.2201 0.2431 0.2430 2.05E+05 2.57E+07
ReplaceMe 0.41 0.6170 0.4430 0.0976 0.3742 0.2747 0.6040 0.3158 0.2638 0.3738 2.30E+02 1.76E+03
LLM-Pruner 0.5033 0.2580 0.0151 0.2643 0.2577 0.2810 0.2182 0.2324 0.2537 $\infty$ 5.67E+05
CoSpaDi 0.4 0.6371 0.4138 0.3029 0.3914 0.2662 0.6850 0.3053 0.2538 0.4069 1.84E+02 1.19E+02
ASVD 0.5196 0.2547 0.0000 0.2601 0.2261 0.1980 0.2134 0.2554 0.2409 1.72E+06 1.72E+07
ReplaceMe 0.5724 0.3635 0.0367 0.2976 0.2543 0.3810 0.2785 0.2300 0.3017 6.99E+02 5.03E+04
LLM-Pruner 0.5038 0.2623 0.0043 0.2630 0.2637 0.2560 0.2172 0.2371 0.2509 1.24E+03 1.23E+06
CoSpaDi 0.5 0.5642 0.3102 0.0390 0.2988 0.2218 0.3770 0.2287 0.2301 0.2837 8.10E+02 7.64E+03
A.7 Cross-Layer Scenario
In Table 6, we report the dictionary size ($k$), sparsity level ($s$), and input/output weight dimensions for each linear transform under a fixed $k/s$ ratio. This includes the query, key, value, and output projections within the attention block, as well as the up, down, and gated projections in the gated MLP block. The dictionary and sparsity dimensions are adjusted according to the target compression rate.
Table 6: Compression configurations for Llama2 7B weight matrices across varying compression rates (20%–50%). Each row specifies sparsity parameters ($k$, $s$, $k/s$ ratio) and weight dimensions ($d_{1}$, $d_{2}$) for different weight types (Query, Key, Value, Up, Gate, Down, Out), grouped by compression rate and group size (1 or 2).
Weight Type Compression Rate Group size Dictionary, $k$ Sparsity, $s$ $k/s$ ratio d1 d2
Query 20% 2 3276 1638 2 4096 8192
Key 3276 1638 2 4096 8192
Value 3276 1638 2 4096 8192
Up 4776 2388 2 4096 22016
Gate 4776 2388 2 4096 22016
Down 1 2762 1381 2 11008 4096
Out 2184 1092 2 4096 4096
Query 30% 2 2866 1433 2 4096 8192
Key 2866 1433 2 4096 8192
Value 2866 1433 2 4096 8192
Up 4178 2089 2 4096 22016
Gate 4178 2089 2 4096 22016
Down 1 2416 1208 2 11008 4096
Out 1910 955 2 4096 4096
Query 40% 2 2456 1228 2 4096 8192
Key 2456 1228 2 4096 8192
Value 2456 1228 2 4096 8192
Up 3582 1791 2 4096 22016
Gate 3582 1791 2 4096 22016
Down 1 2072 1036 2 11008 4096
Out 1638 819 2 4096 4096
Query 50% 2 2048 1024 2 4096 8192
Key 2048 1024 2 4096 8192
Value 2048 1024 2 4096 8192
Up 2984 1492 2 4096 22016
Gate 2984 1492 2 4096 22016
Down 1 1726 863 2 11008 4096
Out 1364 682 2 4096 4096
A.8 K-SVD and Power Iteration Analysis
@@
A.9 Use of Large Language Models
The authors acknowledge the use of a large language model (LLM) solely for language editing and grammatical refinement of the current manuscript. All scientific content, analysis, and interpretations presented herein are the sole responsibility of the authors.

### v3

@@ A.1 Derivation of the optimal pair of basis and coefficient matrices
Let us consider the weight matrix $\mathbf{W}$ as a collection of $d_{1}$-dimensional vectors $\mathbf{W}=\begin{bmatrix}\mathbf{w}_{1},&\ldots,\mathbf{w}_{d_{2}}\end{bmatrix}$, where $\bm{w}_{j}\in\mathbb{R}^{d_{1}}$, with $j=1,\ldots,d_{2}$. We seek to approximate each vector $\mathbf{w}_{j}$ as a linear combination of basis vectors spanning a lower-dimensional subspace of $\mathbb{R}^{d_{1}}$. The optimal basis and coefficients can be found by minimizing the total approximation error:
To solve this problem, we first reformulate the objective $\mathcal{J}$ into an equivalent form:
@@ A.5 Low-rank and CoSpaDi Inference Complexity
With proposed truncation of 2-bits in mantissa we can omit term for storing indices of nonzero elements in $\mathbf{S}$ resulting in corrected compression ratio:
$\hat{\gamma}^{SD}=1-\frac{d_{1}k+sd_{2}}{d_{1}d_{2}}$
The rank for the low-rank decomposition could be estimated from the compression ratio with the following equation:
$r=\frac{(1-\gamma^{LR})d_{1}d_{2}}{d_{1}+d_{2}}$
For sparse dictionary based method we defined $\rho=k/s$ and, thus we can estimate both $k$ and $s$ in the following way:
$k=\frac{(1-\hat{\gamma}^{\mathrm{SD}})\,d_{1}d_{2}}{\,d_{1}+\tfrac{d_{2}}{\rho}\,},\qquad s=\tfrac{k}{\rho}.$
Under the same compression ratio for both SVD and CoSpaDi we obtain exactly the same complexity:
$O_{LR}=Nr(d_{1}+d_{2})=N(1-\gamma^{LR})d_{1}d_{2}$
$O_{SD}=Nd_{1}k+Nsd_{2}=N(d_{1}k+d_{2}\frac{k}{\rho})=Nk(d_{1}+\frac{d_{2}}{\rho})=N\frac{(1-\hat{\gamma}^{\mathrm{SD}})\,d_{1}d_{2}}{\,d_{1}+\tfrac{d_{2}}{\rho}\,}(d_{1}+\frac{d_{2}}{\rho})=N(1-\hat{\gamma}^{\mathrm{SD}})\,d_{1}d_{2}$
@@
While, theoretical complexity is the same, in practice inference time depends on the sparsity structure ($K_{\text{active}}$), indexing overhead, and kernel efficiency, as can be observed from Figs. 5, 6 and   7.
@@
A.6 More Benchmarks
We additionally evaluate CoSpaDi and SVD-LLM on the Qwen model family using a set of more recent reasoning-oriented benchmarks. Table 6 reports results for Qwen3-8B and Qwen3-14B across several compression ratios. Overall, CoSpaDi matches or improves upon the SVD-LLM baseline on most benchmarks and settings, indicating that sparse dictionary learning provides a stronger compression–quality trade-off than low-rank factorization beyond the standard evaluation suite.
Table 6: Performance comparison of CoSpaDi vs sota SVD-LLM on Qwen3-8B and Qwen3-14B at different compression levels on different benchmarks. Best results are highlighted with bold.
Model CR IFEval BBH MATH GPQA MUSR MMLU-Pro
Qwen 3 8B 39.2 60.9 52.6 36.2 43.1 47.7
SVDLLM 25.5 41.0 1.1 28.4 39.8 26.3
CoSpaDi 0.2 28.9 45.3 2.0 28.6 42.1 31.5
SVDLLM 22.9 34.4 1.0 25.6 41.4 18.8
CoSpaDi 0.3 25.2 38.2 1.0 24.8 38.4 22.8
SVDLLM 22.7 30.2 0.8 23.1 37.7 11.6
CoSpaDi 0.4 26.1 32.7 0.8 26.0 38.1 16.7
Qwen 3 14B 43.3 63.2 53.3 38.6 40.7 53.3
SVDLLM 25.5 50.1 1.6 29.2 41.5 32.7
CoSpaDi 0.2 29.6 54.2 4.2 30.6 40.7 38.9
SVDLLM 23.0 34.3 1.1 26.4 39.7 12.8
CoSpaDi 0.4 26.6 38.3 1.2 25.5 38.0 15.7
A.7 K-SVD and Power Iteration Analysis